hartnn (hartnn):

A little bit difficult than previous : what is the missing number in sequence: 9,91,__,855.

hartnn (hartnn):

@Hero

OpenStudy (lgbasallote):

200

hartnn (hartnn):

if u want one more number, after 855 comes 1729 this should ring a bell.

hartnn (hartnn):

nopes,try again lg

mathslover (mathslover):

yes... 341

mathslover (mathslover):

Right?

hartnn (hartnn):

341 indeed

hartnn (hartnn):

post your logic after 5 minutes.

OpenStudy (anonymous):

:)

mathslover (mathslover):

I have to go now... can I nw?

mathslover (mathslover):

Or should I mail you and you post from my behalf? hartnn

OpenStudy (lgbasallote):

oh lol

mathslover (mathslover):

"messaged"

OpenStudy (lgbasallote):

odd numbers

OpenStudy (lgbasallote):

right @hartnn ?

mathslover (mathslover):

no

hartnn (hartnn):

any odd numbers? u see the pattern?

mathslover (mathslover):

will u post from my behalf ? hartnn ?

OpenStudy (lgbasallote):

well...i didnt mean it was the pattern

OpenStudy (lgbasallote):

im stating part of it...

hartnn (hartnn):

hmm.. so u are yet to figure out why 341

hartnn (hartnn):

and its totally math

ganeshie8 (ganeshie8):

9 = 1^3+2^3 91 = 3^3+4^3 341 = 5^3+6^3

hartnn (hartnn):

u goth it ganeshie

OpenStudy (mathmate):

The best hint is from 1729, the taxi number.

OpenStudy (lgbasallote):

the lgbariddle master is losing...

hartnn (hartnn):

not at all!

OpenStudy (mathmate):

I have miscalculated 7^3+8^3 as 755 so missed it. :(

hartnn (hartnn):

don't worry @mathmate i will give ample of such opportunities in future

OpenStudy (mathmate):

Thanks! :)